OpenStudy (anonymous):

If you observed a car traveling at a constant rate around a circular track of 100 m diameter in 5 seconds, what is it's rate of motion in m/s? Then what would it be in miles/hour?

OpenStudy (anonymous):

A circle of diameter d=100m has a radius of d/2=50m As \[v=\omega R\] and \[\omega=2\pi f\] \[f=1/T\] The period of the motion is the time taken for one circuit: T=5s So. \[v=\omega R =2\pi fR= 2\pi R/T = (100\pi)/5 = 62.8ms ^{-1}\] which is 153 miles per hour. pretty fast.

OpenStudy (anonymous):

Thank-you, but I have one question. How come you didn't multiply the '2'? This is very helpful, and I am getting the same answer, but I am confused as to why the 2 isn't multiplied.

OpenStudy (anonymous):

The radius which is 100/2=50 but it is then multiplied by 2 again so goes back to 100. so 2R=100. does that help?

OpenStudy (anonymous):

Yes, thank-you very much. You've been very helpful.